The return on shares of Valley Transporter is predicted under the following various economic conditions:
Recession  -0.13
Normal +0.09
Boom +0.22
If each economy state has the same probability of occurring, what is the variance of the stock
Answer #1

Expected Return = (-0.13+0.09+0.22)/3 = 0.06

Variance = (1/3)*(-0.13-0.06)2 + (1/3)*(0.09-0.06)2 +(1/3)*(0.22-0.06)2 = 0.02087 or 2.09%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well